The price of a 3in by 5in poster is $6.50. What is the price of a 6in by 10 in poster?
Answer by richwmiller(17219)

Don't know. You should ask at Kinko's
Your problem doesn't really give enough info.
But we can make a bunch of assumptions.
3 in by 5 in is not a poster but an index card same as a picture postcard.
Perhaps you mean feet and not inches.
3*5=15 square inches
6*10=60 square inches which is 4 times the postcard
So it should cost 4 times as much.
6 by 10 is just a large postcard not a poster either.
4*6.50=$26.00 for a large postcard